# OneForecast: A Universal Framework for Global and Regional Weather Forecasting

### v1

@@ Abstract
Accurate weather forecasts are important for disaster prevention, agricultural planning, and water resource management. Traditional numerical weather prediction (NWP) methods offer physically interpretable high-accuracy predictions but are computationally expensive and fail to fully leverage rapidly growing historical data. In recent years, deep learning methods have made significant progress in weather forecasting, but challenges remain, such as balancing global and regional high-resolution forecasts, excessive smoothing in extreme event predictions, and insufficient dynamic system modeling. To address these issues, this paper proposes a global-regional nested weather forecasting framework based on graph neural networks (GNNs). By combining a dynamic system perspective with multi-grid theory, we construct a multi-scale graph structure and densify the target region to capture local high-frequency features. We introduce an adaptive information propagation mechanism, using dynamic gating units to deeply integrate node and edge features for more accurate extreme event forecasting. For high-resolution regional forecasts, we propose a neural nested grid method to mitigate boundary information loss. Experimental results show that the proposed method performs excellently across global to regional scales and short-term to long-term forecasts, especially in extreme event predictions (e.g., typhoons), significantly improving forecast accuracy. Our codes are available at https://github.com/YuanGao-YG/OneForecast.
@@ 1 Introduction
Figure 1: Forecast results of extreme cyclones. (a) OneForecast’s predicted wind speed for Typhoon Molva (2020) at 850 hPa pressure level with a 60-hour lead time. (b)–(c) the predicted cyclone tracks of Typhoon Yagi (2018) and Typhoon Molva (2020) using different models
Recent developments in deep learning (DL) methods offer new perspectives for weather forecasting. Early spatio-temporal prediction algorithms (Wu et al., 2024a), such as ConvLSTM (Shi et al., 2015) and PredRNN (Wang et al., 2022), focus on regional precipitation. Recent large-scale scientific computing models, like Pangu-weather (Bi et al., 2023), GraphCast (Lam et al., 2023), and NowcastNet (Zhang et al., 2023), achieve significant results in medium- and short-term forecasts and show high potential in extreme event prediction (e.g., precipitation and Typhoon track prediction) (Chen et al., 2024; Espeholt et al., 2022; Gong et al., 2024; Wu et al., 2024b). However, pure AI methods still face several core challenges:
@@ 2.1 Earth-specific Region Refined Graph Encoder
where ${\mathcal{V}}^{G}$ represents the set of lat-lon grid nodes, with a total of $N=H\times W$ nodes; $\mathcal{V}$ represents the mesh nodes. $\mathcal{E}^{(l)}$ denotes the edge set at the $l$-th scale, corresponding to grids of different granularities, where $l=1,2,\dots,L$, and $\mathcal{E}^{(R)}$ represents regional refined edges. $\mathcal{E}^{\mathrm{G}2\mathrm{M}}$ and $\mathcal{E}^{\mathrm{M}2\mathrm{G}}$ are the unidirectional edges that connect lat-lon grid nodes and mesh nodes. All scales share the same set of nodes $\mathcal{V}$. More details can be found in Appendix D.
@@ 2.2 Multi-stream Messaging
where, $\mathbf{W}_{2}\in\mathbb{R}^{3HD\times h}$ is the weight matrix of the linear transformation of the second layer, and $D$ is the feature dimension for each gate. For each head $h$, the gating values $\mathbf{g}_{i}^{(h,e)}$, $\mathbf{g}_{i}^{(h,s)}$, and $\mathbf{g}_{i}^{(h,d)}$ are vectors of dimension $D$, corresponding to the edge feature gate, source node feature gate, and destination node feature gate, respectively.
Subsequently, we use Edge Sum MLP (ESMLP) to perform linear transformation and nonlinear activation on the edge features to generate the updated edge features:
@@
Multi-head Node Attention Mechanism. Compared to traditional message passing mechanisms, multi-head attention mechanisms can more precisely capture complex dependencies between nodes and dynamically adjust the way information is aggregated through attention weights. For each edge $e_{i}=(j\rightarrow k)$, we use an MLP to calculate the attention score:
$\mathbf{a}_{i}=\operatorname{MLP}_{a}\left(\mathbf{e}_{i}^{\text{new}}\right)\in\mathbb{R}^{H},$ (9)
@@
$\mathbf{m}_{k}^{(h)}=\sum_{e_{i}\in\mathcal{E}(k)}\alpha_{i}^{(h)}\cdot\mathbf{e}_{i}^{\text{new}}\in\mathbb{R}^{D_{e}^{\prime}},$ (11)
@@
Finally, we concatenate the aggregated edge features with the original node features and, through an MLP, perform a nonlinear transformation to generate the updated node features:
@@ 2.4 Downstream Tasks
Regional High-Resolution Forecasting. To enhance the accuracy of high-resolution forecasts in specific regions, we propose a neural nested grid method, illustrated in Figure 2(c). This method combines global low-resolution future forecasts with regional high-resolution data to produce detailed forecasts for the target region. We first input the global low-resolution data at time $t$ into the pre-trained global model to obtain the global forecasts $\hat{\mathbf{Z}}_{t+1}^{\text{global}}$ at time $t+1$. We extract $\hat{\mathbf{Z}}_{t+1}^{\text{global1}}$ from $\hat{\mathbf{Z}}_{t+1}^{\text{global}}$, which shares the same spatial range as the region, and $\hat{\mathbf{Z}}_{t+1}^{\text{global2}}$, which includes the boundary of the region (the boundary are defined as two grid points around the region). Both $\hat{\mathbf{Z}}_{t+1}^{\text{global1}}$ and $\hat{\mathbf{Z}}_{t+1}^{\text{global2}}$ are then interpolated to match the resolution of the high-resolution regional data, which are concatenate as $\hat{\mathbf{Z}}_{t+1}^{\text{global}}$ to acted as global force. We then combine the regional high-resolution data at time $t$ with the $\hat{\mathbf{Z}}_{t+1}^{\text{global}}$ to form the input to the regional model. The global forecasts provide the necessary boundary conditions for the regional forecasts. The regional model then produces the high-resolution forecasts for the regional state at time $t+1$:
@@
Long-Term and Esemble Weather Forecasting. The initial condition of the atmospheric state is uncertain, so reasonable quantification of this uncertainty is conducive to improve to forecast performance. In this work, we set N=50. To account for the uncertainty in the atmospheric initial state for long-term ensemble forecasting, we generate $N$ perturbed initial conditions $\mathbf{Z}_{0}^{(n)}$ by adding Perlin noise $\varepsilon^{(n)}$ to $\mathbf{Z}_{0}$ (Chen et al., 2023b). Each perturbed initial condition is input into the model, and through recursive rollout over $T$ time steps, we obtain individual forecasts $\hat{\mathbf{Z}}_{t+1}^{(n)}$. Finally, at each time step $t+1$, we compute the ensemble mean prediction $\hat{\mathbf{Z}}_{t+1}^{\text{ensemble}}=\frac{1}{N}\sum_{n=1}^{N}\hat{\mathbf{Z}}_{t+1}^{(n)}$ by averaging the forecasts from all $N$ ensemble members.
@@ 3 Experiments
Table 1: In all benchmark tests, we compare the performance of our study with 4 baselines. RMSE and ACC are recorded. A small RMSE ($\downarrow$) and a bigger ACC ($\uparrow$) indicate better performance. The best results are in bold, and the second best are with underline.
@@
Pangu-weather (Bi et al., 2023) 0.0921 0.9958 0.1635 0.9868 0.3723 0.9305 0.5576 0.8414 0.6680 0.7763
Fengwu (Chen et al., 2023a) 0.1428 0.9913 0.2584 0.9738 0.5186 0.8958 0.7874 0.8061 1.2370 0.7409
Graphcast (Lam et al., 2023) 0.0887 0.9962 0.1634 0.9878 0.3962 0.9308 0.6026 0.8362 0.7374 0.7625
Fuxi (Chen et al., 2023b) 0.1106 0.9941 0.1838 0.9838 0.4329 0.9055 0.6056 0.8116 0.6913 0.7571
OneForecast(ours) 0.0683 0.9976 0.1294 0.9917 0.2954 0.9577 0.4823 0.8839 0.6306 0.8025
OneForecast(Promotion) 19.11% 0.11% 19.38% 0.38% 20.66% 2.89% 13.49% 5.05% 5.59% 3.38%
@@ 3.1 Benchmarks and Baselines
We conduct the experiments on the weatherbench2 (Rasp et al., 2024) benchmark, a subset of the fifth generation of ECMWF Reanalysis Data (ERA5) (Hersbach et al., 2020). The subset we use includes years from 1959 to 2020, which is 1959-2017 for training, 2018-2019 for validing, and 2020 for testing. We use 5 pressure level variables (each with 13 pressure levels), geopotential (Z), specific humidity (Q), temperature (T), U and V components of wind speed (U and V), and 4 surface level variables 10-meter U and V components of wind (U10M and V10M), 2-meter temperature (T2M), and Mean sea-level pressure (MSLP). For the global weather forecasts, we choose the 1.5° (121 × 240 for the global data) version of weatherbench2 as our dataset. For convenience, we just choose 120 × 240 data to train models. For regional weather forecasts, we use the original 0.25° (721 × 240 for the global data) ERA5 data. More details can be found in the Appendix B. We compare the forecast performance of OneForecast with 4 data-driven models (Pangu (Bi et al., 2023), Grapcast (Lam et al., 2023), Fengwu (Chen et al., 2023a), and Fuxi (Chen et al., 2023b)). And these models are trained using same framework and settings. We also compare extreme event prediction with IFS HRES, the state-of-the-art numerical method.
@@ 3.2 Comparison with state-of-the-art methods
We utilize two metrics, RMSE (Root Mean Square Error) and ACC (Anomalous Correlation Coefficient), to evaluate the forecast performance. More details can be found in E.1. Since the magnitudes of different variables vary greatly, we first normalize the 69 variables and then calculate the indicators for the 1300 initial conditions. As shown in Table 1, OneForecast achieves satisfactory performance compared with the state-of-the-art models. As shown in Figure 3, OneForecast are closer with the ground truth. We also show the forecast results of several important variables in Figure 4, which are not normalized. Our OneForecast performs better than other models. This improvement is primarily attributed to integrating the proposed adaptive message passing module, which enhances OneForecast’s ability to model the relationships between atmospheric states across different regions of the earth that allows for the simulation of atmospheric dynamics at various spatial and temporal scales adaptively. In summary, the forecasts exhibit more consistency with the actual physical field, effectively mitigate over-smoothing, and demonstrate superior predictive performance, particularly for extreme atmospheric values.
@@ 3.3 Regional High Resolution Forecast
Although the training cost of low-resolution forecast models is relatively low, their prediction results lack sufficient details. However, directly training on high-resolution regional data often results in poor regional forecasts performance due to issues such as missing boundary conditions and limited data samples. Although the regional forecasts method proposed by  (Oskarsson et al., 2024) improves prediction accuracy, it remains constrained by the absence of global information in high-resolution regional models, leaving room for improvement in forecast results. In contrast, the Neural Nesting Grid (NNG) method proposed in this study incorporates boundary conditions and global future information, enabling more accurate high-resolution regional predictions. Furthermore, NNG makes full use of the forecast results of global models, which achieves high-resolution regional forecasts at an exceptionally low training cost. Therefore, as shown in Figure 6, we only demonstratively conduct high-resolution predictions for two regional variables without requiring training on all variables (e.g., the 69 variables used for training global models). It can be seen that the poor long-term inference results are poor when only using regional data for training. Graph-EFM takes into account boundary conditions and the effect is improved. And our proposed NNG not only takes into account regional boundary conditions, but also makes full use of the future forecast information of the global model, which achieves stable long-term forecast performance.
@@ 3.4 Extreme Events Assessment
Extreme events, such as tropical cyclones, can significantly impact human society. In this section, we evaluate our model’s ability in forecasting those extreme cyclones. As shown in Figure 1, OneForecast achieves competitive performance in typhoon tracking during two extreme events, Yagi (2018) and Molva (2020). The details of tracking algorithm can be found in Appendix E.3. Additionally, we download the forecast results of baseline models (e.g., Pangu, Fuxi and Graphcast) from weatherbench2, which is trained using high resolution (0.25°) data, to better illustrate the performance of the baselines. Although OneForecast uses lower-resolution (1.5°) data, which may limit its capacity to predict cyclones, it nevertheless shows strong forecast skills in tracking tropical cyclones comparing with the baselines.
Table 2: Comparison results between deterministic forecasts and uncertain ensemble forecast (ENS), the best results are in bold. We select 10 initial conditions, and for ENS, each results contains 50 members.
Model Metrics
4-day 10-day
RMSE ACC RMSE ACC
Pangu 0.3721 0.9334 0.6778 0.7862
Pangu (ENS) 0.3488 0.9402 0.5640 0.8236
Fengwu 0.5309 0.8913 1.1019 0.7454
Fengwu (ENS) 0.5107 0.8971 0.9288 0.7781
Graphcast 0.6897 0.7443 0.8591 0.6724
Graphcast (ENS) 0.3805 0.9378 0.6400 0.8101
Fuxi 0.4425 0.9061 0.7364 0.7556
Fuxi (ENS) 0.4053 0.9175 0.6001 0.7999
OneForecast 0.2939 0.9594 0.6153 0.8252
OneForecast (ENS) 0.3221 0.9536 0.5385 0.8446

### v2

@@ Abstract
Accurate weather forecasts are important for disaster prevention, agricultural planning, etc. Traditional numerical weather prediction (NWP) methods offer physically interpretable high-accuracy predictions but are computationally expensive and fail to fully leverage rapidly growing historical data. In recent years, deep learning models have made significant progress in weather forecasting, but challenges remain, such as balancing global and regional high-resolution forecasts, excessive smoothing in extreme event predictions, and insufficient dynamic system modeling. To address these issues, this paper proposes a global-regional nested weather forecasting framework (OneForecast) based on graph neural networks. By combining a dynamic system perspective with multi-grid theory, we construct a multi-scale graph structure and densify the target region to capture local high-frequency features. We introduce an adaptive messaging mechanism, using dynamic gating units to deeply integrate node and edge features for more accurate extreme event forecasting. For high-resolution regional forecasts, we propose a neural nested grid method to mitigate boundary information loss. Experimental results show that OneForecast performs excellently across global to regional scales and short-term to long-term forecasts, especially in extreme event predictions. Codes link: https://github.com/YuanGao-YG/OneForecast.
@@ 1 Introduction
Figure 1: Forecast results of extreme typhoons. (a) OneForecast’s predicted wind speed for Typhoon Molva (2020) at 850 hPa pressure level with a 60-hour lead time. (b)–(c) the predicted cyclone tracks of Typhoon Yagi (2018) and Typhoon Molva (2020) using different models
Recent developments in deep learning (DL) models offer new perspectives for weather forecasting. Early spatio-temporal prediction algorithms (Wu et al., 2024a), such as ConvLSTM (Shi et al., 2015) and PredRNN (Wang et al., 2022), focus on regional precipitation. Recent large-scale scientific computing models, like Pangu-weather (Bi et al., 2023), GraphCast (Lam et al., 2023), and NowcastNet (Zhang et al., 2023), achieve significant results in medium- and short-term forecasts and show high potential in extreme event prediction (e.g., precipitation and Typhoon track prediction) (Chen et al., 2024; Espeholt et al., 2022; Gong et al., 2024; Wu et al., 2024b). However, pure AI methods still face several core challenges:
@@ 2.1 Earth-specific Region Refined Graph Encoder
where ${\mathcal{V}}^{G}$ represents the set of lat-lon grid nodes, with a total of $N=H\times W$ nodes; $\mathcal{V}$ represents the mesh nodes. $\mathcal{E}^{(l)}$ denotes the edge set at the $l$-th scale, corresponding to grids of different granularities, where $l=1,2,\dots,L$, and $\mathcal{E}^{(R)}$ represents regional refined edges. $\mathcal{E}^{\mathrm{G}2\mathrm{M}}$ and $\mathcal{E}^{\mathrm{M}2\mathrm{G}}$ are the unidirectional edges that connect lat-lon grid nodes and mesh nodes. All scales share the same set of nodes $\mathcal{V}$. More details can be found in Appendix E.
@@ 2.2 Multi-stream Messaging
where, $\mathbf{W}_{2}\in\mathbb{R}^{3HD\times h}$ is the weight matrix of the linear transformation of the second layer, and $D$ is the feature dimension for each gate. For each head $h$, the gating values $\mathbf{g}_{i}^{(h,e)}$, $\mathbf{g}_{i}^{(h,s)}$, and $\mathbf{g}_{i}^{(h,d)}$ are vectors of dimension $D$, corresponding to the edge feature gate, source node feature gate, and destination node feature gate, respectively. Subsequently, we use Edge Sum MLP (ESMLP) (Pfaff et al., 2020) to perform linear transformation and nonlinear activation on the edge features to generate the updated edge features:
@@
Multi-head Node Attention Mechanism. Compared to traditional message passing mechanisms, multi-head attention mechanisms can more precisely capture complex dependencies between nodes and dynamically adjust the way information is aggregated through attention weights. For each edge $e_{i}=(j\rightarrow k)$, we use a MLP to calculate the attention score:
$\mathbf{a}_{i}=\operatorname{MLP}_{a}\left(\mathbf{e}_{i}^{\text{new}}\right)\in\mathbb{R}^{H}.$ (9)
@@
$\mathbf{m}_{k}^{(h)}=\sum_{e_{i}\in\mathcal{E}(k)}\alpha_{i}^{(h)}\cdot\mathbf{e}_{i}^{\text{new}}\in\mathbb{R}^{D_{e}^{\prime}}.$ (11)
@@
Finally, we concatenate the aggregated edge features with the original node features and, through a MLP, perform a nonlinear transformation to generate the updated node features:
@@ 2.4 Downstream Tasks
Regional High-Resolution Forecasting. To enhance the accuracy of high-resolution forecasts in specific regions, we propose a neural nested grid method, illustrated in Figure 2(c). This method combines global low-resolution future forecasts with regional high-resolution data to produce detailed forecasts for the target region. We first input the global low-resolution data at time $t$ into the pre-trained global model to obtain the global forecasts $\hat{\mathbf{Z}}_{t+1}^{\text{g}}$ at time $t+1$. We extract $\hat{\mathbf{Z}}_{t+1}^{\text{global1}}$ from $\hat{\mathbf{Z}}_{t+1}^{\text{g}}$, which shares the same spatial range as the region, and $\hat{\mathbf{Z}}_{t+1}^{\text{global2}}$, which includes the boundary of the region (the boundary are defined as two grid points around the region). Both $\hat{\mathbf{Z}}_{t+1}^{\text{global1}}$ and $\hat{\mathbf{Z}}_{t+1}^{\text{global2}}$ are then interpolated to match the resolution of the high-resolution regional data, which are concatenate as $\hat{\mathbf{Z}}_{t+1}^{\text{global}}$ to acted as global forcing. We then combine the regional high-resolution data at time $t$ with the $\hat{\mathbf{Z}}_{t+1}^{\text{global}}$ to form the input of the regional model. The global forecasts provide the necessary boundary conditions for the regional forecasts. The regional model then produces the high-resolution forecasts for the regional state at time $t+1$:
@@
Long-Term and Esemble Weather Forecasting. The initial condition of the atmospheric state is uncertain, so reasonable quantification of this uncertainty is conducive to improve to forecast performance. To account for the uncertainty in the atmospheric initial state for long-term ensemble forecasting, we generate $N$ perturbed initial conditions $\mathbf{Z}_{0}^{(n)}$ by adding Perlin noise $\varepsilon^{(n)}$ to $\mathbf{Z}_{0}$ (Chen et al., 2023b). Each perturbed initial condition is input into the model, and through recursive rollout over $T$ time steps, we obtain individual forecasts $\hat{\mathbf{Z}}_{t+1}^{(n)}$. Finally, at each time step $t+1$, we compute the ensemble mean prediction $\hat{\mathbf{Z}}_{t+1}^{\text{ensemble}}=\frac{1}{N}\sum_{n=1}^{N}\hat{\mathbf{Z}}_{t+1}^{(n)}$ by averaging the forecasts from all $N$ ensemble members. In this work, we set N=50.
@@ 3 Experiments
Table 1: In global weather forecasting task, we compare the performance of our OneForecast with 3 baselines, which are trained in the same framework. The average results for all 69 variables of RMSE (normalized) and ACC are recorded. A small RMSE ($\downarrow$) and a bigger ACC ($\uparrow$) indicate better performance. The best results are in bold, and the second best are with underline.
@@
Pangu-weather (Bi et al., 2023) 0.0826 0.9876 0.1571 0.9581 0.3380 0.8167 0.5092 0.5738 0.6215 0.3542
Graphcast (Lam et al., 2023) 0.0626 0.9928 0.1304 0.9705 0.2861 0.8705 0.4597 0.6692 0.6009 0.4275
Fuxi (Chen et al., 2023b) 0.0987 0.9820 0.1708 0.9511 0.4128 0.7379 0.5972 0.4446 0.6981 0.2391
OneForecast(ours) 0.0549 0.9943 0.1231 0.9737 0.2732 0.8825 0.4468 0.6888 0.5918 0.4457
OneForecast(Promotion) 12.24% 0.15% 5.54% 0.33% 4.50% 1.38% 2.81% 2.92% 1.51% 4.25%
@@ 3.1 Benchmarks and Baselines
We conduct the experiments on the WeatherBench2 (Rasp et al., 2024) benchmark, a subset of the fifth generation of ECMWF Reanalysis Data (ERA5) (Hersbach et al., 2020). The subset we use includes years from 1959 to 2020, which is 1959-2017 for training, 2018-2019 for validating, and 2020 for testing. We use 5 pressure level variables (each with 13 pressure levels), geopotential (Z), specific humidity (Q), temperature (T), U and V components of wind speed (U and V), and 4 surface level variables 10-meter U and V components of wind (U10M and V10M), 2-meter temperature (T2M), and mean sea-level pressure (MSLP). For the global weather forecasting task, we choose the 1.5° (121 × 240 for the global data) version of WeatherBench2 as our dataset. For convenience, we just choose 120 × 240 data to train models. For the regional high-resolution weather forecasting task, we use the original 0.25° (721 × 240 for the global data) ERA5 data. More details can be found in the Appendix C. We conduct 2 types comparison, the first type is the comparison between 1-step supervised models retrained using the same framework and settings, which includes Pangu (Bi et al., 2023), Graphcast (Lam et al., 2023), Fuxi (Chen et al., 2023b), and Ours. The second type is the comparison between the results released by WeatherBench2 (with many finetune tricks), Fengwu (Chen et al., 2023a) (results released by the author), and our finetune model.
@@ 3.2 Comparison with state-of-the-art methods
We utilize four metrics, RMSE, ACC, CSI, and SEDI to evaluate the forecast performance. More details can be found in F.1. Since the magnitudes of different variables vary greatly, we first normalize the 69 variables and then calculate the indicators for the 1300 initial conditions. As shown in Table 1, OneForecast achieves satisfactory performance compared with the state-of-the-art models. As shown in Figure 3, OneForecast are closer with the ground truth. We also show the forecast results of several important variables in Figure 4, which are not normalized. Our OneForecast performs better than other models. More results compared with WeatherBench2 can be found in Appendix G.4. This improvement is primarily attributed to integrating the proposed adaptive message passing module, which enhances OneForecast’s ability to model the relationships between atmospheric states across different regions of the earth that allows for the simulation of atmospheric dynamics at various spatial and temporal scales adaptively. In summary, the forecasts exhibit more consistency with the actual physical field, effectively mitigate over-smoothing, and demonstrate superior predictive performance, particularly for extreme atmospheric values.
@@ 3.3 Regional High Resolution Forecast
Although the training cost of low-resolution forecast models is relatively low, their prediction results lack sufficient details. However, directly training on high-resolution regional data often results in poor regional forecasts performance due to issues such as missing boundary conditions and limited data samples. Although the regional forecasts method proposed by  (Oskarsson et al., 2024) improves prediction accuracy, it remains constrained by the absence of global information in high-resolution regional models. In contrast, our proposed Neural Nesting Grid method (NNG) incorporates boundary conditions and global future information, enabling more accurate high-resolution regional predictions. Furthermore, NNG makes full use of the forecast results of global models, which achieves high-resolution regional forecasts at an exceptionally low training cost. Therefore, as shown in Figure 6, we only demonstratively conduct high-resolution predictions for two regional variables without requiring training on all variables (e.g., the 69 variables used for training global models). It can be seen that the poor long-term inference results are poor when only using regional data for training. Graph-EFM takes into account boundary conditions and the effect is improved. And our proposed NNG not only takes into account regional boundary conditions, but also makes full use of the future forecast information of the global model, which achieves stable long-term forecast performance.
@@ 3.4 Extreme Events Assessment
Extreme events, such as tropical cyclones, can significantly impact human society. In this section, we evaluate our model’s ability in forecasting those extreme cyclones. As shown in Figure 1, OneForecast achieves competitive performance in typhoon tracking during two extreme events, Yagi (2018) and Molva (2020). For Yagi, due to Best Track (Ying et al., 2014) (Lu et al., 2021) doesn’t report intact track, we treat ERA5 as the ground truth. For Molva, we treat Best Track as the ground truth. The details of tracking algorithm can be found in Appendix F.3. Additionally, we download the forecast results of baseline models (e.g., Pangu, Fuxi and Graphcast) from WeatherBench2, which is trained using high resolution (0.25°) data, to better illustrate the performance of the baselines. Although OneForecast uses lower-resolution (1.5°) data, which may limit its capacity to predict cyclones, it nevertheless shows strong forecast skills in tracking tropical cyclones comparing with the baselines.
Table 2: Comparison results of RMSE between deterministic forecasts and ensemble forecasts (ENS), the best results are in bold.
Model Forecast Day
7-day 8-day 9-day 10-day
Pangu 0.4875 0.5321 0.5742 0.6213
Pangu (ENS) 0.4435 0.4743 0.4979 0.5205
Graphcast 0.4440 0.4923 0.5346 0.5823
Graphcast (ENS) 0.4412 0.4759 0.5072 0.5331
Fuxi 0.5928 0.6314 0.6604 0.6968
Fuxi (ENS) 0.4898 0.5175 0.5353 0.5498
OneForecast 0.4268 0.4834 0.5313 0.5809
OneForecast (ENS) 0.4393 0.4699 0.4951 0.5167